An NFL punt covered 41.1 meters and the hang time was 4.44 seconds. At what angle was the football punted?
1
Expert's answer
2019-11-11T15:51:47-0500

The range is


"R=\frac{v^2}{g}\sin{2\theta}"

Time of flight:


"T=\frac{2v}{g}\sin{\theta}\to v=\frac{gT}{2\sin{\theta}}"

"R=\frac{(\frac{gT}{2\sin{\theta}})^2}{g}\sin{2\theta}"

"R=\frac{gT^2}{4}\cot{\theta}"

"41.1=\frac{(9.8)(4.44)^2}{4}\cot{\theta}"

"\cot{\theta}=0.85096"

"\theta =\arcctg{0.85096}=49.6\degree"